OpenStudy (anonymous):

@isaiah97 At the city museum, child admission is $5.70 and adult admission is $9.40. On Tuesday, 176 tickets were sold for a total sales of $1362.10. How many child tickets were sold that day?

OpenStudy (anonymous):

238

OpenStudy (anonymous):

first try wrong

OpenStudy (anonymous):

does it say how many adult tickets where sold

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wtf idk dis college work

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is dat a yeah

OpenStudy (anonymous):

+5.70x−9.401769.40x =1362.10 +−3.70x1654.40 =1362.10 −3.70x =−292.30 x =−292.30−3.70 x =79

OpenStudy (texaschic101):

a + c = 176....and since we are looking for c, rearrange to a = 176 - c 5.70c + 9.40a = 1362.10 now it is just a substitution problem... Let me set it up for you... 5.70c + 9.40(176 - c) = 1362.10 can you do the math and finish....solve for c and you will have your answer

OpenStudy (anonymous):

i did it was 79 ty

OpenStudy (texaschic101):

I was just explaining....not just giving the answer

OpenStudy (anonymous):

i understand it was my question and a friend wanted to help.. i wrote out how i recieved the answer so he would know.. thank you for not just posting answer

OpenStudy (texaschic101):

its all good :)

OpenStudy (anonymous):

@isaiah97 ready for another

OpenStudy (anonymous):

sure u make me fell dumb

OpenStudy (anonymous):

I was close @Missiey

OpenStudy (anonymous):

Felipe's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Felipe $4.15 per pound, and type B coffee costs $5.20 per pound. This month's blend used three times as many pounds of type B coffee as type A, for a total cost of $711.00 . How many pounds of type A coffee were used?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

close 36

OpenStudy (anonymous):

here lets do something different

OpenStudy (anonymous):

freakin genius status yeah I am so smart

OpenStudy (anonymous):

Solve the following system of equations. -7x+6y=6 -4x-9y=-9 x= y=

OpenStudy (anonymous):

-1x+12y 5x-y